OpenStudy (anonymous):

Two objects are moving along separate linear paths where each path is described by postion, d, and time, t. The variable d is measured in meters and the variable t is measured in seconds. The equation describing the graph of the postion of the first object with respect to time is d=2.5t+2.2. The graph of the the position of the second object is a parallel line passing through (t=0,d=1). What is the equation of the second graph. d=2.5t+1 d=-0.4t+1 d=2.5t+3.2

OpenStudy (anonymous):

it would have to have the same slope since it is parallel

OpenStudy (anonymous):

d=-0.4t+1

OpenStudy (anonymous):

and it passes through the point (1,0)

OpenStudy (anonymous):

no dude it has the same slope

OpenStudy (anonymous):

it would be 2.5

OpenStudy (anonymous):

not -0.4

OpenStudy (unklerhaukus):

the equation of a line in slope intercept form is y=mx+b, where m is the slope & b is the y-intercept (the value when x is zero)

OpenStudy (unklerhaukus):

in this equation d is like y & t is like x

OpenStudy (unklerhaukus):

@binks can you tell me the slope of d=2.5t+2.2 ?

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

im confused how you find the equation

OpenStudy (unklerhaukus):

yeah 2.5 is the right slope, parallel lines have the same slope so this rules out one of the options

OpenStudy (unklerhaukus):

the other piece of information in the question says that when t=0, d=1 so if we plug in these d=2.5t+b becomes 1=2.5(0)+b solve this for b

OpenStudy (anonymous):

b=1

OpenStudy (unklerhaukus):

yeah so you have found the slope to be m=2.5 & the y-intercept to be b=1 d=2.5t+1

OpenStudy (unklerhaukus):

understand?

OpenStudy (anonymous):

thank you!!!